Connect English with Mathematics and Graphing Lines
4. Translate each sentence into an algebraic equation.
a) One sixth of a number, increased by 15, is 42.
b) Three times a number, decreased by four, is five more than six times the number.
c) When tickets to a soccer game cost $4 each, the revenue at the gate is $320.
d) the total length of the base and height of a triangle 15cm.
PLeaseeeeeeeeeeeeeeeeeee thankssssssssssssssss alottttttttttttttttttt

Found 2 solutions by eperette, stanbon:
Answer by eperette(173)

a) (1/6)x+15=42
b) 3x-4= 6x+5
c) 4x=320
d) b + h = 15
I do have to tell you that question c and d sound incomplete....

Answer by stanbon(75887)

Connect English with Mathematics and Graphing Lines
4. Translate each sentence into an algebraic equation.

a) One sixth of a number, increased by 15, is 42.
(1/6)x +15 = 42
----------------------------
b) Three times a number, decreased by four, is five more than six times the number.
3x - 4 = 6x + 5
----------------------------
c) When tickets to a soccer game cost $4 each, the revenue at the gate is $320.
5x = 320
-------------------------------
d) the total length of the base and height of a triangle 15cm.
b + h = 15
===========================================
Cheers,
Stan H.